A scientist measures the standard enthalpy change for the following reaction to be -112.0 kJ :

2NO(g) + O2(g) ------> 2NO2(g)

Based on this value and the standard enthalpies of formation for the other substances, the standard enthalpy of formation of NO2(g) is _________ kJ/mol.
1
Expert's answer
2019-04-08T07:55:59-0400
"\Delta H_f(NO)=90.25 \frac{kJ}{mol}"

"\Delta H_{rxn}= \sum \Delta H_f^0(products)-\sum \Delta H_f^0(reactants)""\Delta H_{rxn}= 2\times \Delta H_f^0(NO_2)-(2\times \Delta H_f^0(NO)+1\times \Delta H_f^0(O_2))"

"-112.0 \frac{kJ}{mol}= 2\times\Delta H_f^0(NO_2)-(2\times 90.25 \frac{kJ}{mol}+1\times 0\frac{kJ}{mol})"

"\Delta H_f^0(NO_2)= 34.25 \frac{kJ}{mol}"